OpenStudy (anonymous):

PLEASE HELP ASAP Naming properties: 16.)x-14=5 17.)x+54=-23 18.)5x=-20 19.)x/4=4 20.)3x+11=20 21.)x/3-4=11

OpenStudy (mathstudent55):

Usually to name a property, you show an expression and a second expression in which something was changed, and you need to say which property allows you to go from the the first expression to the second one. I don't know what you want to do here.

OpenStudy (radar):

16. x - 14 = 5 x -14 +14 = 5 + 14 x = 19 If equals are added to equals, the results are equal.

OpenStudy (radar):

I too am not sure what this problem is wanting!!

OpenStudy (anonymous):

its not asking for a result its asking u to name the property,. identity. associative.closure. the property names

OpenStudy (radar):

Sorry,but I have to read too much into that.

OpenStudy (mathstudent55):

You mean the property you need to use to solve each equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hey radar r u good at equations with fractions? story problems? two-step equaions? 1 step equations?

OpenStudy (radar):

Fair to middlin.

OpenStudy (anonymous):

can u answer the question I posted??????

OpenStudy (anonymous):

????????????????????????

OpenStudy (radar):

let me see if I can find it.

OpenStudy (anonymous):

thanks. APPRECIATE you help (: please hurry tho cuz I have bunches of questions

OpenStudy (radar):

I am 74 years old, and hurry was yesterday. Where is the problem posted?

OpenStudy (anonymous):

your 74!!!???????

OpenStudy (radar):

Yes, and I am still looking for your new problem, or are you talking about those at the beginning of this thread.

OpenStudy (anonymous):

name the properties 16.) x-14=5 17.)x+54= -23 18.) 5x= -20 19.)x/4=4 20.) 3x+11=20 21.) x/3-4=11

OpenStudy (anonymous):

same problems.you don't look 74 in your pic jeez lol

OpenStudy (radar):

This is the one I don't understand like for number 17. x + 54 = -23 Equals can be subtracted from equals and the differences are equal This is one of the fundamental rules of algebra. x + 54 - 54 = -23 - 54 x = -77

OpenStudy (radar):

You have already stated that is not what this is about, so I am at a lost.

OpenStudy (anonymous):

): I could try to post a pic like out of the book explaining it???

OpenStudy (radar):

That might help.

OpenStudy (anonymous):

camera not working sorry

OpenStudy (radar):

elp is here. Do you understand what this problem is about @nimcompoop?

OpenStudy (nincompoop):

16.)x-14=5 (axiom) the existence of subtraction 17.)x+54=-23 18.)5x=-20 identity property of multiplication 19.)x/4=4 (axiom) existence of division 20.)3x+11=20 21.)x/3-4=11

OpenStudy (anonymous):

18 looks like the only one close to being what the teacher wants

OpenStudy (anonymous):

only its just identity

OpenStudy (radar):

This problem is too vague and ambigious for me. Good bye and good luck.

OpenStudy (anonymous):

ok

OpenStudy (mary.rojas):

@Psymon ?

OpenStudy (anonymous):

what?????

OpenStudy (mary.rojas):

I am asking him to help you lol

OpenStudy (mary.rojas):

:( he cant help, he is leaving now..... hold on

OpenStudy (anonymous):

ok )':

OpenStudy (anonymous):

I dont know how can they be properities! all of them are equations

OpenStudy (anonymous):

unless you mean the property to solve the equations..

OpenStudy (anonymous):

name the properties you would use to solve equation. associative multiplication. closure. identity. distributive.

OpenStudy (nincompoop):

Property Example 1. Commutative Property of Addition a + b = b + a 2 + 3 = 3 + 2 2. Commutative Property of Multiplication a • b = b • a 2 • ( 3 ) = 3 • ( 2 ) 3. Associative Property of Addition a + ( b + c ) = ( a + b ) + c 2 + ( 3 + 4 ) = ( 2 + 3 ) + 4 4. Associative Property of Multiplication a • ( b • c ) = ( a • b ) • c 2 • ( 3 • 4 ) = ( 2 • 3 ) • 4 5. Distributive Property a • ( b + c ) = a • b + a • c 2 • ( 3 + 4 ) = 2 • 3 + 2 • 4 6. Additive Identity Property a + 0 = a 3 + 0 = 3 7. Multiplicative Identity Property a • 1 = a 3 • 1 = 3 8. Additive Inverse Property a + ( -a ) = 0 3 + (-3) = 0 9. Multiplicative Inverse Property Note: a cannot = 0 10. Zero Property a • 0 = 0 5 • 0 = 0 A set is closed (under an operation) if and only if the operation on two elements of the set produces another element of the set. If an element outside the set is produced, then the operation is not closed.

OpenStudy (anonymous):

that doesn't help me with my questions tho

OpenStudy (amistre64):

lol ... oh but it does

OpenStudy (amistre64):

im assuming you need to solve the equations and define the property you used in each step

OpenStudy (nincompoop):

Axiom (Existence of Subtraction). For every a in R, there is an element in R denoted by −a such that a + (−a) = 0. Note. a − b is an abbreviation of a + (−b).

OpenStudy (anonymous):

yeah,it is the right answers but I have no clue which one matches with what problem lol and what is axiom?

OpenStudy (amistre64):

its not a match up, its a solve while defining which one you use ... or at least thats what i get from it

OpenStudy (anonymous):

no u don't solve. just define property

OpenStudy (nincompoop):

20.)3x+11=20 closure property - the 3 in 3x denotes an odd and 11 is an odd; after solving you get an odd number 9. therefore it is closure property

OpenStudy (amistre64):

some of your problems do not have a onestep solution property to apply

OpenStudy (raden):

i think it is just about open sentences --"

OpenStudy (anonymous):

does anyone know the answers??? u just name property? I have ALOT more questions, so please noonee leave

OpenStudy (nincompoop):

I've been giving you the answers LOL

OpenStudy (anonymous):

not understanding then

OpenStudy (anonymous):

can u write the numbers wih the answers next to problem?

OpenStudy (nincompoop):

16.)x-14=5 (axiom) existence of subtraction; identity property of addition x - 14 + 14 = 5 + 14 x + 0 = 19 x = 19

OpenStudy (anonymous):

thanks. I don't know what the rest of them are tho

OpenStudy (nincompoop):

REMEMBER THIS? 6. Additive Identity Property and this example? a + 0 = a 3 + 0 = 3

OpenStudy (nincompoop):

you have to apply yourself on the definition of properties…

OpenStudy (anonymous):

yeah but idk which number it goes with I took geometry last year. this is alg 2 work. the question I asked had geometry mixed in.. idk which number gets what property?????????????

OpenStudy (anonymous):

nincompoop do u know how to help me sole story problems? equations with fractions??

OpenStudy (anonymous):

also after u help me with property questions pleaseee

OpenStudy (nincompoop):

I can guide you

OpenStudy (anonymous):

well im still lost on 16-21

OpenStudy (nincompoop):

but create a new thread

OpenStudy (anonymous):

what u mean?

OpenStudy (anonymous):

would 17 be additive inverse property??

OpenStudy (nincompoop):

go all the way up of this question you created then click close the question. now create a new question on the "ask a question" on the upper left hand corner

OpenStudy (nincompoop):

but the key to answering this problem is what @amistre64 indicated solve them and then observe which properties were used or are applicable from beginning to the end.

OpenStudy (anonymous):

I waited for 2 hours for help with prop questions, please don't leave

OpenStudy (nincompoop):

hehehe well close this and ask a new question